# Intrinsic training dynamics of deep neural networks

### v1

@@ Conclusion
hence $U^{\prime}u$ is colinear with $u$. Since this holds for any choice of $u$, we deduce indeed that $U^{\prime}$ is proportional to $\mathrm{Id}_{n}$. A similar reasoning yields that $V^{\prime}\propto\mathrm{Id}_{m}$. This concludes the proof of (25).
@@
$\Delta_{S}(2U^{\top}U-S)+(2U^{\top}U-S)\Delta_{S}=[\Delta_{A},S].$
with $\Delta_{S}$ (resp. $\Delta_{A}$) the symmetric (resp. skew symmetric) part of $\Delta$ (so that $\Delta=\Delta_{S}+\Delta_{A}$). Denote $\mathcal{S}_{r}$ (resp. $\mathcal{A}_{r}$) the set of $r\times r$ symmetric (resp. skew symmetric) matrices and $L$ the Lyapunov operator defined by:
@@
As $S\neq\lambda\mathrm{Id}_{r}$, by Lemma˜I.1 there exists a skew-symmetric matrix $\Delta_{A}\in\mathcal{A}_{r}$ such that $\Delta_{A},S]\neq 0$. As $U^{\top}U+V^{\top}V$ is positive definite (as either $U$ or $V$ has full column-rank) its eigenvalues $\lambda_{i}>0$ satisfy $\lambda_{i}+\lambda_{j}\neq 0$, so (see e.g. Bartels & Stewart (1972)) in particular the Lyapunov operator: $L:\mathcal{S}_{r}\to\mathcal{S}_{r}$ is invertible. Since $\Delta_{A},S]=\Delta_{A}S-S\Delta_{A}\in\mathcal{S}_{r}$, we obtain that there exists $\Delta_{S}\neq 0$ such that $L(\Delta_{S})=[\Delta_{A},S]$. This particular choice of $\Delta_{S}$ and $\Delta_{A}$ exhibits a parameter $\theta^{\prime}=(U\Delta,-V\Delta^{\top})$ that satifies $\theta^{\prime}\in\mathcal{V}\subseteq\mathrm{ker}\partial\phi(\theta)\cap\mathrm{ker}\partial\mathbf{h}(\theta)$. We now show that $\theta^{\prime}\notin\mathrm{ker}\partial{M}(\theta)$. We proceed by contradiction: if $\theta^{\prime}\in\mathrm{ker}\partial{M}(\theta)$ then, by (25), there exists $\mu\in\mathbb{R}$ such that $U(\Delta^{\top}+\Delta)U^{\top}=\mu\mathrm{Id}_{n}\quad\text{and}\quad V(\Delta^{\top}+\Delta)V^{\top}=-\mu\mathrm{Id}_{m}$ that is to say
@@
When $r\leq\max(m,n)$ and since $U$, $V$ are full rank, at least one of the two matrices $U$ or $V$ is full column rank $r$. Without loss of generality, let us assume that $U$ is full column rank. Then $U^{\top}U$ is invertible and we deduce that,
@@
Moreover if (as we indeed show below) $\mathrm{range}U^{\top}\cap\mathrm{range}V^{\top}\neq\{0\}$, then by considering $z=U^{\top}x=V^{\top}y\neq 0$ for some $x,y\in\mathbb{R}^{r}$, one deduces from (26) that $\mu\|x\|_{2}^{2}=2x^{\top}U\Delta_{S}U^{\top}x=z^{\top}\Delta_{S}z=2y^{\top}V\Delta_{S}V^{\top}y=-\mu\|y\|_{2}^{2}$ and thus $\mu=0$. Hence $\Delta_{S}=0$ by (27), contradicting $L(\Delta_{S})=[\Delta_{A},S]\neq 0$, which shows that $\theta^{\prime}\notin\mathrm{ker}\partial{M}(\theta)$.
@@
As both $U_{t=0}$ and $V_{t=0}$ have full rank it remains the case in a neighborhood $\Omega$ of $\theta_{0}$. Moreover as $r\leq\text{max}(n,m)$ then one of the two matrices has a full column rank on $\Omega$. In particular the vertical concatenation $\begin{pmatrix}U\\ V\end{pmatrix}$ has full rank (equal to $r$) on $\Omega$ as $r\leq\text{max}(n,m)\leq n+m$.
@@
We now show by contradiction that for any ${\Omega^{\prime}}\subseteq{\Omega}$, $\theta_{0}$ does not satisfy the intrinsic metric on $\Omega^{\prime}$. Let us assume there exists a neighborhood $\Omega^{\prime}\subseteq\Omega$ of $\theta_{0}$ and a set of conservation laws $\mathbf{h}_{0}$ for $\phi$ and a function $K_{\theta_{0}}$ such that $M(\theta)=K_{\theta_{0}}(\phi(\theta))$ for each $\theta\in\mathcal{M}_{\theta_{0}}^{\mathbf{h_{0}}}\cap\Omega^{\prime}$, where $\mathcal{M}_{\theta_{0}}^{\mathbf{h_{0}}}\coloneqq\{\theta:\mathbf{h_{0}}(\theta)=\mathbf{h_{0}}(\theta_{0})\}$. As the family of conservation laws $\mathbf{h}$ is complete on $\Omega$ (and in particular on $\Omega^{\prime}$) and as $\mathrm{Lie}(\mathbb{W}_{\phi})(\theta)$ has a constant dimension on $\Omega$ (and thus on $\Omega^{\prime}$) by (Marcotte et al., 2023, Proposition 4.3), using (Marcotte et al., , Proposition 2.12) yields that $\mathcal{M}_{\theta_{0}}^{\mathbf{h_{0}}}\coloneqq\{\theta:\mathbf{h_{0}}(\theta)=\mathbf{h_{0}}(\theta_{0})\}\supset\mathcal{M}_{\theta_{0}}^{\mathbf{h}}\coloneqq\{\theta:\mathbf{h}(\theta)=\mathbf{h}(\theta_{0})\}$. Thus the function $K_{\theta_{0}}$ also satisfies $M(\theta)=K_{\theta_{0}}(\phi(\theta))$ on $\mathcal{M}_{\theta_{0}}^{\mathbf{h}}$, hence $\mathbf{h}$ satisfies assumption $i)$ of Theorem˜2.14. As the rank of $\partial h(\theta)$ is constant on $\Omega^{\prime}$, we deduce by Theorem˜2.14 the inclusion (7), which contradicts the previous step. ∎
Appendix J Proof of Proposition˜4.5.
See 4.5
@@
Appendix K Proof of Theorem˜4.6.
See 4.6
@@
Let us first outline the main steps of the proof. We first show that the equalities $Z_{L}Z_{L}^{\top}=Q_{0}(U_{L}U_{L}^{\top})$ and $Z_{L}^{\top}Z_{L}=R_{L-1}(U_{1}^{\top}U_{1})$ hold on the whole trajectory. Then we prove that this implies the expression of $S_{j}$ (resp. of $T_{j}$) in terms of $U_{L}U_{L}^{\top}$ (resp. of $U_{1}^{\top}U_{1}$) along the whole trajectory too. Finally we show that along the whole trajectory $U_{L}U_{L}^{\top}$ and $U_{1}^{\top}U_{1}$ (and therefore all $S_{j}$’s and $T_{j}$’s) are entirely characterized by $Z_{L}=\phi_{\mathtt{Lin}}(\theta_{L})$ and the initial conditions (captured by $\boldsymbol{\lambda}$). This will thus imply that $\theta_{L}(0)$ satisfies the intrinsic dynamic property on $\mathbb{R}^{D}$ with respect to $\phi_{\mathtt{Lin}}$.
@@
Since $U_{j+1}^{\top}U_{j+1}-U_{j}U_{j}^{\top}$ is a set of conservation laws for $\phi_{\mathtt{Lin}}$, the fact that the relaxed balanced conditions (13) hold at initialization implies that they hold along the whole trajectory.
@@
where we used (13) for $i=\ell-1$. Denoting $\hat{P}_{\ell-1}(x):=P_{\ell-1}(x-\lambda_{\ell-1})$ we obtain $Z_{\ell}Z_{\ell}^{\top}=U_{\ell}\hat{P}_{\ell-1}(U_{\ell}^{\top}U_{\ell})U_{\ell}^{\top}=U_{\ell}U_{\ell}^{\top}\hat{P}_{\ell-1}(U_{\ell}U_{\ell}^{\top})=P_{\ell}(U_{\ell}U_{\ell}^{\top})$. This concludes the induction.
@@
where $\boldsymbol{\lambda}^{\prime}=(\lambda^{\prime}_{L-j-1},\ldots,\lambda^{\prime}_{1}):=(\lambda_{L-1},\ldots,\lambda_{j+1})$. By the first step, we obtain the desired expression.
Similar computations with $\theta^{\prime}=(U_{1}^{\top},\ldots,U_{j-1}^{\top})$, $Z^{\prime}=U_{1}^{\top}\ldots U_{j-1}^{\top}$ and ${\boldsymbol{\lambda}}^{\prime}=(-\lambda_{1},\ldots,-\lambda_{j-2})$ show the desired expression for $T_{j}=Z^{\prime}Z^{\prime\top}$ and $Z_{L}^{\top}Z_{L}$ as well.
@@
As we show below, a consequence of the relaxed balancedness conditions (13) is that all eigenvalues of the positive semi-definite matrix $U_{L}U_{L}^{\top}$ belong to the interval $I:=[\max(0,a_{1},\ldots,a_{L-1}),\infty)$. Thus, considering any eigenvalue $e\geq 0$ of the positive semi-definite matrix $Z_{L}Z_{L}^{\top}$, it is enough to show that the polynomial equation $R(X):=Q_{0}(X)-e=0$ admits a unique root in this interval.
@@
To conclude the proof, we show that indeed all eigenvalues of $U_{L}U_{L}^{\top}$ belong to $I:=[\max(0,a_{1},\ldots,a_{L-1}),\infty)$. Denote $\sigma_{i}=\inf\text{sp}(U_{i}U_{i}^{\top})$, $1\leq i\leq L$. Since each matrix $U_{\ell}$ is square and positive semi-definite, we have $\text{sp}(U_{i}U_{i}^{\top})=\text{sp}(U_{i}^{\top}U_{i})\subseteq[0,\infty)$ for every $1\leq i\leq L$, and by (13) we also have $\text{sp}(U_{i+1}U_{i+1}^{\top})=\lambda_{i}+\text{sp}(U_{i}U_{i}^{\top})$, hence $\sigma_{i+1}=\sigma_{i}+\lambda_{i}\geq 0$ for $1\leq i\leq L-1$. An easy recursion shows that $\sigma_{i}\geq\max(0,\sum_{j=1}^{i-1}\lambda_{j})$ for $1\leq i\leq L$, hence the result. ∎
We now anticipate a slight generalization part of the results of Theorem˜4.6 that will be used later in the proof of Theorem˜4.8.
@@
Step 3: Conclusion. Applying the integral formula of Step 1with the bound from Step 2 gives
@@
for every $A,\Delta$ with $A,A+\Delta\in B_{R(h)}$, which is (32). ∎
@@
We also observe that by the definition of $\eta$ in (30), since $h=1/L$, we have for each $1\leq\ell\leq N$
@@
Induction step. Assume (34) holds at rank $\ell-1$. Since $Z_{\ell}=U^{\prime}_{\ell}Z_{\ell-1}$,
@@
Write $Q_{\ell-1}(x):=P_{\ell-1}(x-b_{\ell-1})$. From (33) and the definition of $M_{\ell-1}=U^{\prime}_{\ell-1}[U^{\prime}_{\ell-1}]^{\top}$ one gets
@@
By the definition of $C_{1}$, the matrices $A=U^{\prime}_{\ell}{}^{\top}U^{\prime}$, $\Delta=-r_{\ell-1}^{\prime}$, satisfy $\max(\|A\|,\|\Delta\|)\leq\|A\|+\|\Delta\|\leq C_{U}^{2}+h^{2}\eta\leq 1+C_{1}h$. Moreover, with the same induction that has led to (28), the polynomial $P_{\ell-1}(x)$ has all its roots bounded by $Lh^{2}C_{\lambda}$, hence $Q_{\ell-1}(x):=P_{\ell-1}(x-b_{\ell-1})$ has all its roots bounded by $(L+1)h^{2}C_{\lambda}\leq 2C_{\lambda}h=C_{0}h$, therefore we can apply Lemma˜K.2 to obtain, with $K=K(C_{0},C_{1})=C_{0}e^{C_{0}(1+C_{1})}+e^{C_{1}}$:
@@
Step 2: Factorisation of $P_{N}$. With the same induction that has led to (28), we have
@@
Applying (34) with $\ell=N$ and recalling $S_{j}=Z_{N}Z_{N}^{\top}$ yields
@@
Since $M_{N}=U^{\prime}_{N}{U^{\prime}_{N}}^{\top}=U_{L-1}{U_{L-1}}^{\top}$, we recover (31) as claimed. ∎
Appendix L Proof of Proposition˜4.7.
See 4.7
@@
For convenience we recall the state equation (17) for $Z_{s}$, where $s\in[0,1]$ indicates depth:
@@
Let $\theta:[t\in[0,T]\mapsto\theta(t)\in\mathcal{X}]\in\mathcal{C}^{1}([0,T],\mathcal{X})$ be the solution of the gradient flow given by the family of coupled ODE (16)
@@
To compute the gradient $\mathfrak{g}_{s}=\frac{\partial\ell}{\partial\mathcal{A}_{s}}$, we introduce the adjoint variable (Pontryagin et al. (1962)) $\Lambda_{s}(t)$, which satisfies the adjoint equation
@@
By differentiating (40) with respect to $s$, we get:
@@
Then by using the adjoint equation (39) and the state equation (35), one has
@@
Step 4: Conclusion. By substituting the computed expressions into (L), one obtains as claimed that
@@
Step 3: symmetry of the mixed derivatives. Evaluating the identity (45) above pointwise in $s$ and writing $\mathcal{A}(t,s)=[\theta(t)](s)$ gives
@@
L.2 We now show (40).
More precisely, to show (40), we will both prove that
@@
which will indeed give (40).
We briefly explain why, for a given $t$, the matrix $Z_{s}(t)$ never loses its invertibility when $s\in[0,1]$ varies, by showing that the determinant can never reach 0. As

### v2

@@ F.2.2 Rank properties
hence $U^{\prime}u$ is colinear with $u$. Since this holds for any choice of $u$, we deduce indeed that $U^{\prime}$ is proportional to $\mathrm{Id}_{n}$. A similar reasoning yields that $V^{\prime}\propto\mathrm{Id}_{m}$. This concludes the proof of equation 27.
@@
$\Delta_{S}(2U^{\top}U-S)+(2U^{\top}U-S)\Delta_{S}=[\Delta_{A},S],$
with $\Delta_{S}$ (resp. $\Delta_{A}$) the symmetric (resp. skew-symmetric) part of $\Delta$ (so that $\Delta=\Delta_{S}+\Delta_{A}$). Denote $\mathcal{S}_{r}$ (resp. $\mathcal{A}_{r}$) the set of $r\times r$ symmetric (resp. skew-symmetric) matrices and $L$ the Lyapunov operator defined by:
@@
As $S\neq\lambda\mathrm{Id}_{r}$, by Lemma˜I.1 there exists a skew-symmetric matrix $\Delta_{A}\in\mathcal{A}_{r}$ such that $\Delta_{A},S]\neq 0$. As $U^{\top}U+V^{\top}V$ is positive definite (as either $U$ or $V$ has full column-rank) its eigenvalues $\lambda_{i}>0$ satisfy $\lambda_{i}+\lambda_{j}\neq 0$, so (see e.g. Bartels and Stewart (1972)) in particular the Lyapunov operator: $L:\mathcal{S}_{r}\to\mathcal{S}_{r}$ is invertible. Since $\Delta_{A},S]=\Delta_{A}S-S\Delta_{A}\in\mathcal{S}_{r}$, we obtain that there exists $\Delta_{S}\neq 0$ such that $L(\Delta_{S})=[\Delta_{A},S]$. This particular choice of $\Delta_{S}$ and $\Delta_{A}$ exhibits a parameter $\theta^{\prime}=(U\Delta,-V\Delta^{\top})$ that satisfies $\theta^{\prime}\in\mathcal{V}\subseteq\mathrm{ker}\partial\phi(\theta)\cap\mathrm{ker}\partial\mathbf{h}(\theta)$. We now show that $\theta^{\prime}\notin\mathrm{ker}\partial{M}(\theta)$. We proceed by contradiction: if $\theta^{\prime}\in\mathrm{ker}\partial{M}(\theta)$ then, by equation 27, there exists $\mu\in\mathbb{R}$ such that $U(\Delta^{\top}+\Delta)U^{\top}=\mu\mathrm{Id}_{n}\quad\text{and}\quad V(\Delta^{\top}+\Delta)V^{\top}=-\mu\mathrm{Id}_{m}$ that is to say
@@
When $r\leq\max(m,n)$ and since $U$, $V$ are full rank, at least one of the two matrices $U$ or $V$ is full column rank $r$. Without loss of generality let us assume that $U$ is full column rank. Then $U^{\top}U$ is invertible and we deduce that,
@@
Moreover if (as we indeed show below) $\mathrm{range}(U^{\top})\cap\mathrm{range}(V^{\top})\neq\{0\}$, then by considering $z=U^{\top}x=V^{\top}y\neq 0$ for some $x\in\mathbb{R}^{n}$ and $y\in\mathbb{R}^{m}$, one deduces from equation 28 that $\mu\|x\|_{2}^{2}=2x^{\top}U\Delta_{S}U^{\top}x=z^{\top}\Delta_{S}z=2y^{\top}V\Delta_{S}V^{\top}y=-\mu\|y\|_{2}^{2}$ and thus $\mu=0$. Hence $\Delta_{S}=0$ by equation 29, contradicting $L(\Delta_{S})=[\Delta_{A},S]\neq 0$, which shows that $\theta^{\prime}\notin\mathrm{ker}\partial{M}(\theta)$.
@@
As both $U_{t=0}$ and $V_{t=0}$ have full rank it remains the case in a neighborhood $\Omega$ of $\theta_{0}$. Moreover as $r\leq\text{max}(n,m)$ then one of the two matrices has full column rank on $\Omega$. In particular the vertical concatenation $\begin{pmatrix}U\\ V\end{pmatrix}$ has full rank (equal to $r$) on $\Omega$ as $r\leq\text{max}(n,m)\leq n+m$.
@@
We now show by contradiction that for any ${\Omega^{\prime}}\subseteq{\Omega}$, $\theta_{0}$ does not satisfy the intrinsic metric property on $\Omega^{\prime}$. Let us assume there exists a neighborhood $\Omega^{\prime}\subseteq\Omega$ of $\theta_{0}$ and a set of conservation laws $\mathbf{h}_{0}$ for $\phi$ and a function $K_{\theta_{0}}$ such that $M(\theta)=K_{\theta_{0}}(\phi(\theta))$ for each $\theta\in\mathcal{M}_{\theta_{0}}^{\mathbf{h_{0}}}\cap\Omega^{\prime}$, where $\mathcal{M}_{\theta_{0}}^{\mathbf{h_{0}}}\coloneqq\{\theta:\mathbf{h_{0}}(\theta)=\mathbf{h_{0}}(\theta_{0})\}$. As the family of conservation laws $\mathbf{h}$ is complete on $\Omega$ (and in particular on $\Omega^{\prime}$) and as $\mathrm{Lie}(\mathbb{W}_{\phi})(\theta)$ has a constant dimension on $\Omega$ (and thus on $\Omega^{\prime}$) by (Marcotte et al., 2023, Proposition 4.3), using (Marcotte et al., 2025, Proposition 2.12) yields that $\mathcal{M}_{\theta_{0}}^{\mathbf{h_{0}}}\coloneqq\{\theta:\mathbf{h_{0}}(\theta)=\mathbf{h_{0}}(\theta_{0})\}\supset\mathcal{M}_{\theta_{0}}^{\mathbf{h}}\coloneqq\{\theta:\mathbf{h}(\theta)=\mathbf{h}(\theta_{0})\}$. Thus the function $K_{\theta_{0}}$ also satisfies $M(\theta)=K_{\theta_{0}}(\phi(\theta))$ on $\mathcal{M}_{\theta_{0}}^{\mathbf{h}}$, hence $\mathbf{h}$ satisfies assumption $i)$ of Theorem˜2.14. As the rank of $\partial h(\theta)$ is constant on $\Omega^{\prime}$, we deduce by Theorem˜2.14 the inclusion equation (ii), which contradicts the previous step.
B. Finally we show that such a $\theta_{0}$ does not satisfy the strong intrinsic dynamic property either.
This is a direct consequence of the following lemma.
Lemma I.3.
Consider a two-layer linear network parameterized by $\theta=(U,V)$, $\phi(\theta)=\phi_{\mathtt{Lin}}(\theta)=UV^{\top}$, and an initialization $\theta_{0}=(U_{t=0},V_{t=0})$ such that the vertical concatenation $\begin{pmatrix}U_{t=0}\\ V_{t=0}\end{pmatrix}$ has full rank. If $\theta_{0}$ satisfies the strong intrinsic dynamic property with respect to $\phi$ on $\Omega$, then it also satisfies the intrinsic metric property with respect to $\phi$ on some open neighborhood $\Omega^{\prime}$ of $\theta_{0}$.
Proof of Lemma˜I.3.
Consider the function $K_{\theta_{0}}$ which existence is guaranteed by the fact that $\theta_{0}$ satisfies the strong intrinsic dynamic property. We will use the Chow-Rashevskii theorem to show that there is a neighborhood $\Omega^{\prime}\subseteq\Omega$ of $\theta_{0}$ such that $\mathcal{M}_{\theta_{0}}\cap\Omega^{\prime}$ is attainable by patching a finite number of trajectories $\dot{\theta}_{k}(t)=-\nabla w_{k}(\theta_{k}(t))$, each initiated at the ending point of the previous one, defined via fields
$w_{k}(\cdot)\in\mathcal{F}:=\{w(\cdot)=\partial\phi^{\top}(\cdot)\nabla f(\phi(\cdot)):f\in C^{\infty}\}.$
We will further show that there is a piecewise $\mathcal{C}^{2}$ function $f$ such that this is feasible with a continuous trajectory $\theta(t)$ such that $\dot{\theta}(t)=-\nabla\ell(\theta(t))$ with $\ell=f\circ\phi$ for each $t$ such $\theta(t)\in\Omega$ and $f$ is differentiable at $\phi(\theta(t))$. The strong intrinsic dynamic property of $\theta_{0}$ with respect to $\phi$ thus yields $M(\theta(t))=K_{\theta_{0}}(\phi(\theta(t)))$ at every time, and in particular $M(\theta)=K_{\theta_{0}}(\phi(\theta))$ at the end point of the trajectory. This shows that $\theta_{0}$ satisfies the intrinsic metric property with respect to $\phi$ on $\Omega^{\prime}$.
To exploit the Chow-Rashevskii theorem, we first observe that $\mathcal{F}=$-$\mathcal{F}$, and that by standard Lie algebra calculus, since every $w(\cdot)\in\mathcal{F}$ can be written as
$w(\theta)=\sum_{i=1}^{d}a_{i}(\theta)\nabla\phi_{i}(\theta)$
we have $\mathrm{Lie}(\mathcal{F})(\theta)\subseteq\mathrm{Lie}(\mathbb{W}_{\phi})(\theta)$ where $\mathbb{W}_{\phi}\coloneqq\mathrm{span}\{\nabla\phi_{i}(\cdot)\}$. Vice-versa, considering $e_{i}\in\mathbb{R}^{d}$ the $i$-th canonical vector, since $f_{i}(\phi):=\langle e_{i},\phi\rangle$ is $\mathcal{C}^{\infty}$ with $\nabla f(\phi)=e_{i}$, we get $w_{i}(\cdot):=\partial\phi^{\top}(\cdot)\nabla f_{i}(\cdot)=\nabla\phi_{i}(\cdot)\in\mathcal{F}$, hence $\mathrm{Lie}(\mathbb{W}_{\phi})(\theta)\subseteq\mathrm{Lie}(\mathcal{F})(\theta)$. Moreover, as $\begin{pmatrix}U_{t=0}\\ V_{t=0}\end{pmatrix}$ has full rank it remains the case in a neighborhood $\Omega^{\prime\prime}\subset\Omega$ of $\theta_{0}$. With the same arguments as in step 4 of the proof of Theorem˜3.9 above, the vector-valued function $\mathbf{h}$ contains a complete set of conservation laws of $\phi=\phi_{\texttt{Lin}}$, and by (Marcotte et al., 2023, Propositions 4.2 and 4.3) one thus has $\mathrm{Lie}(\mathcal{F})(\theta)=\mathrm{Lie}(\mathbb{W}_{\phi})(\theta)=T_{\theta}(\mathcal{M}_{\theta_{0}}^{\mathbf{h}}\cap\Omega^{\prime})$ (with $T_{\theta}$ the tangent plane at $\theta$) for each $\theta\in\mathcal{M}_{\theta_{0}}^{\mathbf{h}}\cap\Omega^{\prime}$. Finally, choose an open neighborhood $\Omega^{\prime}\subseteq\Omega^{\prime\prime}$ of $\theta_{0}$ such that $\mathcal{M}_{\theta_{0}}^{\mathbf{h}}\cap\Omega^{\prime}$ is a connected set: all the assumptions of the Chow-Rashevskii theorem (Jurdjevic, 1997, Theorem 3) hold on $\Omega^{\prime}$, thus the attainable sets of $\mathcal{F}$ from $\theta_{0}$ on $\Omega^{\prime}$ is exactly $\mathcal{M}_{\theta_{0}}^{\mathbf{h}}\cap\Omega^{\prime}$. This means that for every $\theta\in\mathcal{M}_{\theta_{0}}^{\mathbf{h}}\cap\Omega^{\prime}$ there is a trajectory $\theta(t),t\in[0,T]$ such that $\theta(t)\in\Omega^{\prime}$ at every time, $\theta(0)=\theta_{0}$, $\theta(T)=\theta$, and $0,T]=\cup_{k}[t_{k},t_{k+1})$ with $\dot{\theta}(t)=w_{k}(\theta(t))$ for $t\in(t_{k},t_{k+1})$, with $w_{k}\in\mathcal{F}$. Without loss of generality the trajectory does not self-intersect (if it does at times $\tau<\tau^{\prime}$, we can shorten it by concatenating the trajectories on $0,\tau]$ and $\tau^{\prime},T]$). Since there are $\mathcal{C}^{\infty}$ functions $f_{k}$ such that $w_{k}=\nabla(f_{k}\circ\phi)$, and since the trajectory $\theta(t)$ does not self-intersect, there is a piecewise $\mathcal{C}^{\infty}$ function $f$ that matches each $f_{k}$ on each piece of the trajectory $\theta(\cdot)$, hence $\dot{\theta}(t)=-\nabla\ell(\theta(t))$ with $\ell=f\circ\phi$, for every $t\in(t_{k},t_{k+1})$ and every $k$.
Appendix J Proof of Proposition˜3.10.
See 3.10
@@
Appendix K Proof of Theorem˜3.11.
See 3.11
@@
Let us first outline the main steps of the proof. We first show that the equalities $Z_{L}Z_{L}^{\top}=Q_{0}(U_{L}U_{L}^{\top})$ and $Z_{L}^{\top}Z_{L}=R_{L-1}(U_{1}^{\top}U_{1})$ hold on the whole trajectory. Then we prove that this implies the expression of $S_{j}$ (resp. of $T_{j}$) in terms of $U_{L}U_{L}^{\top}$ (resp. of $U_{1}^{\top}U_{1}$) along the whole trajectory too. Finally we show that along the whole trajectory $U_{L}U_{L}^{\top}$ and $U_{1}^{\top}U_{1}$ (and therefore all $S_{j}$’s and $T_{j}$’s) are entirely characterized by $Z_{L}=\phi_{\mathtt{Lin}}(\theta_{L})$ and the initial conditions (captured by $\bm{\lambda}$). This will thus imply that $\theta_{L}(0)$ satisfies the intrinsic dynamic property on $\mathbb{R}^{D}$ with respect to $\phi_{\mathtt{Lin}}$.
@@
Since $U_{j+1}^{\top}U_{j+1}-U_{j}U_{j}^{\top}$ is a set of conservation laws for $\phi_{\mathtt{Lin}}$, the fact that the relaxed balanced conditions equation 12 hold at initialization implies that they hold along the whole trajectory.
@@
where we used equation 12 for $i=\ell-1$. Denoting $\hat{P}_{\ell-1}(x):=P_{\ell-1}(x-\lambda_{\ell-1})$ we obtain $Z_{\ell}Z_{\ell}^{\top}=U_{\ell}\hat{P}_{\ell-1}(U_{\ell}^{\top}U_{\ell})U_{\ell}^{\top}=U_{\ell}U_{\ell}^{\top}\hat{P}_{\ell-1}(U_{\ell}U_{\ell}^{\top})=P_{\ell}(U_{\ell}U_{\ell}^{\top})$. This concludes the induction.
@@
where $\bm{\lambda}^{\prime}=(\lambda^{\prime}_{L-j-1},\ldots,\lambda^{\prime}_{1}):=(\lambda_{L-1},\ldots,\lambda_{j+1})$. By the first step we obtain the desired expression.
Similar computations with $\theta^{\prime}=(U_{1}^{\top},\ldots,U_{j-1}^{\top})$, $Z^{\prime}=U_{1}^{\top}\ldots U_{j-1}^{\top}$ and ${\bm{\lambda}}^{\prime}=(-\lambda_{1},\ldots,-\lambda_{j-2})$ show the desired expression for $T_{j}=Z^{\prime}Z^{\prime\top}$ and $Z_{L}^{\top}Z_{L}$ as well.
@@
As we show below, a consequence of the relaxed balancedness conditions equation 12 is that all eigenvalues of the positive semi-definite matrix $U_{L}U_{L}^{\top}$ belong to the interval $I:=[\max(0,a_{1},\ldots,a_{L-1}),\infty)$. Thus, considering any eigenvalue $e\geq 0$ of the positive semi-definite matrix $Z_{L}Z_{L}^{\top}$, it is enough to show that the polynomial equation $R(X):=Q_{0}(X)-e=0$ admits a unique root in this interval.
@@
To conclude the proof, we show that indeed all eigenvalues of $U_{L}U_{L}^{\top}$ belong to $I:=[\max(0,a_{1},\ldots,a_{L-1}),\infty)$. Denote $\sigma_{i}=\inf\text{sp}(U_{i}U_{i}^{\top})$, $1\leq i\leq L$. Since each matrix $U_{\ell}$ is square and positive semi-definite, we have $\text{sp}(U_{i}U_{i}^{\top})=\text{sp}(U_{i}^{\top}U_{i})\subseteq[0,\infty)$ for every $1\leq i\leq L$, and by equation 12 we also have $\text{sp}(U_{i+1}U_{i+1}^{\top})=\lambda_{i}+\text{sp}(U_{i}U_{i}^{\top})$, hence $\sigma_{i+1}=\sigma_{i}+\lambda_{i}\geq 0$ for $1\leq i\leq L-1$. An easy recursion shows that $\sigma_{i}\geq\max(0,\sum_{j=1}^{i-1}\lambda_{j})$ for $1\leq i\leq L$, hence the result. ∎
We now anticipate a slight generalization part of the results of Theorem˜3.11 that will be used later in the proof of Theorem˜3.13.
@@
Step 3: Conclusion. Applying the integral formula for Step 1 with the bound from Step 2 gives
@@
for every $A,\Delta$ with $A,A+\Delta\in B_{R(h)}$, which is equation 34. ∎
@@
We also observe that by the definition of $\eta$ in equation 32, since $h=1/L$, we have for each $1\leq\ell\leq N$
@@
Induction step. Assume equation 36 holds at rank $\ell-1$. Since $Z_{\ell}=U^{\prime}_{\ell}Z_{\ell-1}$,
@@
Write $Q_{\ell-1}(x):=P_{\ell-1}(x-b_{\ell-1})$. From equation 35 and the definition of $M_{\ell-1}=U^{\prime}_{\ell-1}[U^{\prime}_{\ell-1}]^{\top}$ one gets
@@
By the definition of $C_{1}$, the matrices $A=U^{\prime}_{\ell}{}^{\top}U^{\prime}$, $\Delta=-r_{\ell-1}^{\prime}$, satisfy $\max(\|A\|,\|\Delta\|)\leq\|A\|+\|\Delta\|\leq C_{U}^{2}+h^{2}\eta\leq 1+C_{1}h$. Moreover, with the same induction that has led to equation 30, the polynomial $P_{\ell-1}(x)$ has all its roots bounded by $Lh^{2}C_{\lambda}$, hence $Q_{\ell-1}(x):=P_{\ell-1}(x-b_{\ell-1})$ has all its roots bounded by $(L+1)h^{2}C_{\lambda}\leq 2C_{\lambda}h=C_{0}h$, therefore we can apply Lemma˜K.2 to obtain, with $K=K(C_{0},C_{1})=C_{0}e^{C_{0}(1+C_{1})}+e^{C_{1}}$:
@@
Step 2: Factorisation of $P_{N}$. With the same induction that has led to equation 30, we have
@@
Applying equation 36 with $\ell=N$ and recalling $S_{j}=Z_{N}Z_{N}^{\top}$ yields
@@
Since $M_{N}=U^{\prime}_{N}{U^{\prime}_{N}}^{\top}=U_{L-1}{U_{L-1}}^{\top}$, we recover equation 33 as claimed. ∎
Appendix L Proof of Proposition˜3.12.
See 3.12
@@
For convenience we recall the state equation equation 16 for $Z_{s}$, where $s\in[0,1]$ indicates depth:
@@
Let $\theta:[t\in[0,T]\mapsto\theta(t)\in\mathcal{X}]\in\mathcal{C}^{1}([0,T],\mathcal{X})$ be the solution of the gradient flow given by the family of coupled ODE equation 15
@@
To compute the gradient $\mathfrak{g}_{s}=\frac{\partial\ell}{\partial\mathcal{A}_{s}}$, we introduce the adjoint variable (Pontryagin et al., 1962) $\Lambda_{s}(t)$, which satisfies the adjoint equation
@@
By differentiating equation 42 with respect to $s$, we get:
@@
Then by using the adjoint equation equation 41 and the state equation equation 37, one has
@@
Step 4: Conclusion. By substituting the computed expressions into equation L, one obtains as claimed that
@@
Step 3: symmetry of the mixed derivatives. Evaluating the identity equation 47 above pointwise in $s$ and writing $\mathcal{A}(t,s)=[\theta(t)](s)$ gives
@@
L.2 We now show equation 42.
More precisely, to show equation 42, we will both prove that
@@
which will indeed give equation 42.
We briefly explain why for a given $t$ the matrix $Z_{s}(t)$ never loses its invertibility when $s\in[0,1]$ varies, by showing that the determinant can never reach 0. As